karina147:

A researcher collected the heights of male seniors at a high school. The data collected are normally distributed with a mean of 70.21 inches and a standard deviation of 1.35. What is the z-score of a high school senior who is 68 inches tall? Enter your answer rounded to the nearest hundredth, like this: 4.23

Vocaloid:

to calculate the z-score of some value x z-score = (x - x̄)/σ, where x is the value being examined, x̄ is the mean, σ is standard deviation so in your case, x = 68, x̄ = the mean = 70.21, σ = standard deviation = 1.35 plug in the appropriate values, calculate the z-score, and round to the nearest hundredth (2 digits)

karina147:

Thanks but whats the answer?

Vocaloid:

The purpose of this website is to teach, rather than to give away answers. Can you try plugging the values into the formula I gave you?

karina147:

Okay thank you

karina147:

Do i multiply them together?

Vocaloid:

the formula I gave you: z-score = (x - x̄)/σ your x-value is 68, your x̄ is 70.21, your σ is 1.35 plugging them in: (68 - 70.21) / (1.35) = ?

karina147:

-2.9835?

Vocaloid:

order of operations. 68 - 70.21 = -2.21 -2.21 divided by 1.35 ---> -1.64

karina147:

-3.277

Vocaloid:

(68 - 70.21) / (1.35) = -1.64 = your solution

karina147:

1.34

Vocaloid:

this is the z-score formula z-score = (x - x̄)/σ you want the z-score of a high schooler who is 68 inches tall, so x = 68 the mean x̄ is 70.21 inches the standard deviation σ is 1.35 when you plug these into the formula, you get (68 - 70.21) / (1.35) = -1.64, that's it, that's your solution.

karina147:

ooh -1.64 is my solution? Thank you

karina147:

Nope its the wrong answer